OpenStudy (anonymous):

I am really confused! Solve the system by substitution. -x - y - z = -8 -4x + 4y +5z = 7 2x + 2z = 4

OpenStudy (anonymous):

Step 1: Take the first equation and solve for a variable. x= -y-z+8 Step 2: Plug the new equation into a different equation and simplify. [[I will plug it into the 3rd equation]] 2(-y-z+8)+2z=4 -2y-2z+16+2z=4 -2y+16=4 -2y=-12 y=6

OpenStudy (anonymous):

Step 3: Plug in x and y into the last equation. -4((-6)-z+8)+4(6)+5z=7 24+4z-32+24+5z=7 16+9z=7 9z=-9 z=-1

OpenStudy (anonymous):

Step 4: Plug in y and z into the original equation -x-(6)-(-1)=-8 -x-6+1=-8 -x-5=-8 -x=-3 x=3

OpenStudy (anonymous):

SO x=3, y=6, and z=-1

OpenStudy (anonymous):

Thank you so much for breaking it down for me. That makes a lot of sense!